# Combating Confirmation Bias: A Unified Pseudo-Labeling Framework for Entity Alignment

### v3

@@ Abstract
Entity alignment (EA) aims at identifying equivalent entity pairs across different knowledge graphs (KGs) that refer to the same real-world identity. It has been a compelling but challenging task that requires the integration of heterogeneous information from different KGs to expand the knowledge coverage and enhance inference abilities. To circumvent the shortage of prior seed alignments provided for training, recent EA models utilize pseudo-labeling strategies to iteratively add unaligned entity pairs predicted with high confidence to the seed alignments for model training. However, the adverse impact of confirmation bias during pseudo-labeling has been largely overlooked, thus hindering entity alignment performance. To systematically combat confirmation bias, we propose a new Unified Pseudo-Labeling framework for Entity Alignment (UPL-EA) that explicitly alleviates pseudo-labeling errors to boost the performance of entity alignment. UPL-EA achieves this goal through two key innovations: (1) Optimal Transport (OT)-based pseudo-labeling uses discrete OT modeling as an effective means to determine entity correspondences and reduce erroneous matches across two KGs. An effective criterion is derived to infer pseudo-labeled alignments that satisfy one-to-one correspondences; (2) Parallel pseudo-label ensembling refines pseudo-labeled alignments by combining predictions over multiple models independently trained in parallel. The ensembled pseudo-labeled alignments are thereafter used to augment seed alignments to reinforce subsequent model training for alignment inference. The effectiveness of UPL-EA in eliminating pseudo-labeling errors is both theoretically supported and experimentally validated. Our extensive results and in-depth analyses demonstrate the superiority of UPL-EA over 15 competitive baselines and its utility as a general pseudo-labeling framework for entity alignment.
@@ 1 Introduction
Over the last decade, there has been a surge of research efforts dedicated to entity alignment across KGs. Most mainstream EA models are embedding-based; they embed KGs into a shared latent embedding space so that similarities between entities can be measured via their embeddings for alignment inference. To leverage structural information in KGs, more recent EA models exploit the power of Graph Neural Networks (GNNs) to encode KG structures for entity alignment. Methods like GCN-Align [52] utilize GCNs to learn better entity embeddings by aggregating features from neighboring entities. However, GCNs and their variants suffer from an over-smoothing issue [24, 13], where the embeddings of entities among local neighborhoods become indistinguishably similar as the number of convolution layers increases. To alleviate over-smoothing during GCN neighborhood aggregation, recent works [54, 53, 63] use a highway strategy [35] on GCN layers, which “mixes” the smoothed entity embeddings with the original features. Despite achieving competitive results, these methods require an abundant amount of pre-aligned entity pairs (known as prior seed alignments) provided for training, which are labor-intensive and costly to acquire in real-world KGs. To tackle the shortage of prior seed alignments, recently proposed models, such as BootEA [40], IPTransE [62], MRAEA [21], and RNM [63], adopt a bootstrapping strategy that iteratively selects unaligned entity pairs predicted with high confidence as pseudo-labeled alignments and adds them to seed alignments for subsequent model training. The bootstrapping strategy, originating from the field of statistics, is also referred to as pseudo-labeling—a predominant learning paradigm proposed to tackle label scarcity in semi-supervised learning.
@@ 2.1 Problem Statement
Formally, given two KGs, $\mathcal{G}_{1}=\{\mathcal{E}_{1},\mathcal{R}_{1},\mathcal{T}_{1}\}$ and $\mathcal{G}_{2}=\{\mathcal{E}_{2},\mathcal{R}_{2},\mathcal{T}_{2}\}$, the task of entity alignment (EA) aims to discover a set of one-to-one equivalent entity pairs =$\mathcal{I}=\{(e_{i},e_{j})\in\mathcal{E}_{1}\times\mathcal{E}_{2}|\;e_{i}\equiv e_{j}\}$ between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, where $e_{1}\in\mathcal{E}_{1}$, $e_{2}\in\mathcal{E}_{2}$, and $\equiv$ indicates an equivalence relationship between $e_{1}$ and $e_{2}$. In many cases, a small set of equivalent entity pairs $\mathcal{S}\subset\mathcal{I}$, known as prior seed alignments, is initially provided for training. Apart from the entities included in $\mathcal{S}$, there are two sets of unaligned entities $\mathcal{E}^{U}_{1}\subset\mathcal{E}_{1}$ and $\mathcal{E}^{U}_{2}\subset\mathcal{E}_{2}$ in $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively.
In this work, we address real-world scenarios where prior seed alignments $\mathcal{S}$ are often scarce due to high labeling costs. We focus on the task of pseudo-labeling-based entity alignment, which aims to leverage both prior seed alignments and unaligned entity pairs to more effectively train an EA model in a transductive semi-supervised setting. This is achieved by selecting a set of unaligned entity pairs as pseudo-labeled alignments $\mathcal{S}_{t}\subset\mathcal{E}^{U}_{1}\times\mathcal{E}^{U}_{2}$ in each iteration $t$, and using $\mathcal{S}_{t}$ to iteratively augment seed alignments $\mathcal{S}$, i.e., $\mathcal{S}\leftarrow\mathcal{S}\cup\mathcal{S}_{t}$, for subsequent model training.
@@ 3.2 Unified Pseudo-Labeling Strategy
After training the EA model, the learned entity embeddings can be used for alignment inference. However, as prior seed alignments initially provided for training are often limited, the performance of entity alignment can be suboptimal. Therefore, pseudo-labeling strategies can be designed to select a set of unaligned entity pairs as pseudo-labeled alignments, which are used to augment seed alignments for boosting model training. However, as previously discussed in Section 2.2, a naive strategy inevitably introduces a considerable number of pseudo-labeling errors, leading to confirmation bias. To address this issue, we propose UPL, a unified pseudo-labeling framework that explicitly aims to mitigate conflicted misalignments and one-to-one misalignments.
@@ 3.2.1 OT-based Pseudo-Labeling
$T^{*}=\mathop{\arg\min}\limits_{T}\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\bm{C}_{e_{i},T(e_{i})}.$ (4)
A critical aspect of the above objective is defining a reliable measure of the transport cost. One might directly use the distances between the learned entity embeddings. However, when only a limited number of prior seed alignments are available for training, the learned entity embeddings can be uninformative, particularly during the early training stages before the EA model has converged. As a result, using these embeddings to calculate the distances for defining the transport cost can be error-prone. To address this, we resort to rectifying the embedding distance using relational neighborhood matching [63], which complements the training of the EA model, especially during the early stages, for learning better entity embeddings and providing a more reliable cost measure for OT modeling. The principle of distance rectification is to leverage relational contexts within local neighborhoods to help determine the extent to which two entities should be aligned. Intuitively, if two entities $e_{i}\in\mathcal{E}_{1}$ and $e_{j}\in\mathcal{E}_{2}$ share more aligned neighboring entities/relations, the distance between their embeddings should be smaller, indicating a higher likelihood of being aligned to each other. Based on this intuition, the transport cost for OT-based pseudo-labeling is defined as follows:
$\bm{C}_{e_{i},e_{j}}={\color[rgb]{0,0,0}d(e_{i},e_{j})-\lambda s(e_{i},e_{j})},\;e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2},$ (5)
@@
$s(e_{i},e_{j})=\frac{\sum_{(r,e_{k})\in\mathcal{M}_{(e_{i},e_{j})}}P_{1}(r,e_{k})P_{2}(r,e_{k})}{|\mathcal{N}_{e}(e_{i})|+|\mathcal{N}_{e}({e}_{j})|},$ (6)
@@
where $\mathcal{N}_{e}(e_{i})$ and $\mathcal{N}_{e}(e_{j})$ denote the sets of neighboring entities for $e_{i}$ and $e_{j}$, respectively. $P_{1}(r,e_{k})$ and $P_{2}(r,e_{k})$ indicate the reciprocal frequency of triplets associated with neighboring tuple $(r,e_{k})$ for $e_{i}$ in $\mathcal{T}_{1}$ and $e_{j}$ in $\mathcal{T}_{2}$, respectively.
@@
The objective in Eq. (4) defines a hard assignment optimization problem, which, however, does not scale well. To enable more efficient optimization and to allow for a more flexible alignment configuration, we reformulate this objective as a discrete OT problem, where the optimal transport plan is considered as a coupling matrix $\bm{P}^{*}\in\mathbb{R}_{+}^{|\mathcal{E}^{U}_{1}|\times|\mathcal{E}^{U}_{2}|}$ between two discrete distributions. Denote $\bm{\mu}$ and $\bm{\nu}$ as two discrete probability distributions over all entities $\{e_{i}|e_{i}\in\mathcal{E}^{U}_{1}\}$ and $\{e_{j}|e_{j}\in\mathcal{E}^{U}_{2}\}$, respectively. Without any alignment preference, the two discrete distributions $\bm{\mu}$ and $\bm{\nu}$ are assumed to follow a uniform distribution such that $\bm{\mu}=\frac{1}{|\mathcal{E}^{U}_{1}|}\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\delta_{e_{i}}$ and $\bm{\nu}=\frac{1}{|\mathcal{E}^{U}_{2}|}\sum_{e_{j}\in\mathcal{E}^{U}_{2}}\delta_{e_{j}}$, where $\delta_{e_{i}}$ and $\delta_{e_{j}}$ are the Dirac function centered on $e_{i}$ and $e_{j}$, respectively. Both $\bm{\mu}$ and $\bm{\nu}$ are bounded to sum up to one: $\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\bm{\mu}(e_{i})=\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\frac{1}{|\mathcal{E}^{U}_{1}|}=1$ and $\sum_{e_{j}\in\mathcal{E}^{U}_{2}}\bm{\mu}(e_{j})=\sum_{e_{j}\in\mathcal{E}^{U}_{2}}\frac{1}{|\mathcal{E}^{U}_{2}|}=1$. Accordingly, the OT objective is formulated to find the optimal coupling matrix $\bm{P}^{*}$ between $\bm{\mu}$ and $\bm{\nu}$:
$\bm{P}^{*}=$ $\mathop{\arg\min}\limits_{\bm{P}\in\Pi(\bm{\mu},\bm{\nu})}\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\sum_{e_{j}\in\mathcal{E}^{U}_{2}}\bm{P}_{e_{i},e_{j}}\cdot\bm{C}_{e_{i},e_{j}},$ (7)
subject to: $\sum_{e_{j}\in\mathcal{E}^{U}_{2}}\bm{P}_{e_{i},e_{j}}=\bm{\mu}(e_{i})=\frac{1}{|\mathcal{E}^{U}_{1}|},$
$\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\bm{P}_{e_{i},e_{j}}=\bm{\nu}(e_{j})=\frac{1}{|\mathcal{E}^{U}_{2}|},$
$\bm{P}_{e_{i},e_{j}}\geq 0,\forall e_{i}\in\mathcal{E}^{U}_{1},\forall e_{j}\in\mathcal{E}^{U}_{2},$
@@
where $\Pi(\bm{\mu},\bm{\nu})=\{\bm{P}\in\mathbb{R}_{+}^{|\mathcal{E}^{U}_{1}|\times|\mathcal{E}^{U}_{2}|}|\;\bm{P}\mathbf{1}_{|\mathcal{E}^{U}_{2}|}=\bm{\mu},\;\bm{P}^{\top}\mathbf{1}_{|\mathcal{E}^{U}_{1}|}=\bm{\nu}\}$ is the set of all joint probability distributions with marginal probabilities $\bm{\mu}$ and $\bm{\nu}$, $\mathbf{1}_{n}$ denotes an $n$-dimensional vector of ones. $\bm{P}$ is a coupling matrix signifying probabilistic alignments between two unaligned entity sets $\mathcal{E}^{U}_{1}$ and $\mathcal{E}^{U}_{2}$. Therefore, $\bm{P}_{e_{i},e_{j}}$ indicates the amount of probability mass transported from $\bm{\mu}(e_{i})$ to $\bm{\nu}(e_{j})$. A larger value of $\bm{P}_{e_{i},e_{j}}$ indicates a higher likelihood of $e_{i}$ and $e_{j}$ being aligned to each other.
@@
$\bm{P}^{*}=\mathop{\arg\min}\limits_{P\in\Pi(\bm{\mu},\bm{\nu})}\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\sum_{e_{j}\in\mathcal{E}^{U}_{2}}\bm{P}_{e_{i},e_{j}}\cdot\bm{C}_{e_{i},e_{j}}+\beta\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\sum_{e_{j}\in\mathcal{E}^{U}_{2}}\bm{P}_{e_{i},e_{j}}\text{log}\bm{P}_{e_{i},e_{j}},$ (8)
@@
${\color[rgb]{0,0,0}\widehat{\mathcal{S}}_{t}}=\{(e_{i},e_{j})|\;\bm{P}^{*}_{e_{i},e_{j}}>\frac{1}{2\cdot\text{min}(|\mathcal{E}^{U}_{1}|,|\mathcal{E}^{U}_{2}|)},\ e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2}\}.$ (9)
@@
Any pseudo-labeled alignment $(e_{i},e_{j}),e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2}$ that satisfies the condition $\bm{P}^{*}_{e_{i},e_{j}}>\frac{1}{2\cdot\text{min}(|\mathcal{E}^{U}_{1}|,|\mathcal{E}^{U}_{2}|)}$ warrants one-to-one correspondence, such that no conflicted entity pairs, $\{(e_{i},e_{k})|\;e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}\}$ and $\{(e_{l},e_{j})|\;e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}\}$, are selected as pseudo-labeled alignments.
@@
Given the optimized coupling matrix $\bm{P}^{*}\in\mathbb{R}_{+}^{|\mathcal{E}^{U}_{1}|\times|\mathcal{E}^{U}_{2}|}$ subject to the constraints of $\sum_{e_{j}\in\mathcal{E}^{U}_{2}}\bm{P}^{*}_{e_{i},e_{j}}=\frac{1}{|\mathcal{E}^{U}_{1}|}$ for all rows ($\forall e_{i}\in\mathcal{E}^{U}_{1}$) and $\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\bm{P}^{*}_{e_{i},e_{j}}=\frac{1}{|\mathcal{E}^{U}_{2}|}$ for all columns ($\forall e_{j}\in\mathcal{E}^{U}_{2}$). Assume that $|\mathcal{E}^{U}_{1}|<|\mathcal{E}^{U}_{2}|$, the decision threshold is $\frac{1}{2\cdot\text{min}(|\mathcal{E}^{U}_{1}|,|\mathcal{E}^{U}_{2}|)}=\frac{1}{2|\mathcal{E}^{U}_{1}|}$. Entity pairs $\{(e_{i},e_{j})|\;\bm{P}^{*}_{e_{i},e_{j}}>\frac{1}{2|\mathcal{E}^{U}_{1}|},e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2}\}$ are selected as pseudo-labeled alignments.
@@
For each pseudo-labeled alignment $(e_{i},e_{j})$ with a probability value $\bm{P}^{*}_{e_{i},e_{j}}>\frac{1}{2|\mathcal{E}^{U}_{1}|}$, we can prove that no conflicted entity pairs $\{(e_{i},e_{k})|\;e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}\}$ associated with $e_{i}$ are selected as pseudo-labeled alignments:
$\bm{P}^{*}_{e_{i},e_{j}}$ $>\frac{1}{2|\mathcal{E}^{U}_{1}|},$
$\sum_{e_{j}\in\mathcal{E}^{U}_{2}}\bm{P}^{*}_{e_{i},e_{j}}-\bm{P}^{*}_{e_{i},e_{j}}$ $<\sum_{e_{j}\in\mathcal{E}^{U}_{2}}\bm{P}^{*}_{e_{i},e_{j}}-\frac{1}{2|\mathcal{E}^{U}_{1}|},$
$\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}\bm{P}^{*}_{e_{i},e_{k}}+\bm{P}^{*}_{e_{i},e_{j}}-\bm{P}^{*}_{e_{i},e_{j}}$ $<\frac{1}{|\mathcal{E}^{U}_{1}|}-\frac{1}{2|\mathcal{E}^{U}_{1}|},$
$\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}\bm{P}^{*}_{e_{i},e_{k}}$ $<\frac{1}{2|\mathcal{E}^{U}_{1}|}.$ (10)

### v4

@@ Abstract
Entity alignment (EA) aims at identifying equivalent entity pairs across different knowledge graphs (KGs) that refer to the same real-world identity. It has been a compelling but challenging task that requires the integration of heterogeneous information from different KGs to expand the knowledge coverage and enhance inference abilities. To circumvent the shortage of seed alignments provided for training, recent EA models utilize pseudo-labeling strategies to iteratively add unaligned entity pairs predicted with high confidence to the seed alignments for model training. However, the adverse impact of confirmation bias during pseudo-labeling has been largely overlooked, thus hindering entity alignment performance. To systematically combat confirmation bias, we propose a new Unified Pseudo-Labeling framework for Entity Alignment (UPL-EA) that explicitly alleviates pseudo-labeling errors to boost the performance of entity alignment. UPL-EA achieves this goal through two key innovations: (1) Optimal Transport (OT)-based pseudo-labeling uses discrete OT modeling as an effective means to determine entity correspondences and reduce erroneous matches across two KGs. An effective criterion is derived to infer pseudo-labeled alignments that satisfy one-to-one correspondences; (2) Parallel pseudo-label ensembling refines pseudo-labeled alignments by combining predictions over multiple models independently trained in parallel. The ensembled pseudo-labeled alignments are thereafter used to augment seed alignments to reinforce subsequent model training for alignment inference. The effectiveness of UPL-EA in eliminating pseudo-labeling errors is both theoretically supported and experimentally validated. Our extensive results and in-depth analyses demonstrate the superiority of UPL-EA over 15 competitive baselines and its utility as a general pseudo-labeling framework for entity alignment.
@@ 1 Introduction
Over the last decade, there has been a surge of research efforts dedicated to entity alignment across KGs. Most mainstream EA models are embedding-based; they embed KGs into a shared latent embedding space so that similarities between entities can be measured via their embeddings for alignment inference. To leverage structural information in KGs, more recent EA models exploit the power of Graph Neural Networks (GNNs) to encode KG structures for entity alignment. Methods like GCN-Align [52] utilize GCNs to learn better entity embeddings by aggregating features from neighboring entities. However, GCNs and their variants suffer from an over-smoothing issue [24, 13], where the embeddings of entities among local neighborhoods become indistinguishably similar as the number of convolution layers increases. To alleviate over-smoothing during GCN neighborhood aggregation, recent works [54, 53, 63] use a highway strategy [35] on GCN layers, which “mixes” the smoothed entity embeddings with the original features. Despite achieving competitive results, these methods require an abundant amount of pre-aligned entity pairs (known as seed alignments) provided for training, which are labor-intensive and costly to acquire in real-world KGs. To tackle the shortage of seed alignments, recently proposed models, such as BootEA [40], IPTransE [62], MRAEA [21], and RNM [63], adopt a bootstrapping strategy that iteratively selects unaligned entity pairs predicted with high confidence as pseudo-labeled alignments and adds them to seed alignments for subsequent model training. The bootstrapping strategy, originating from the field of statistics, is also referred to as pseudo-labeling—a predominant learning paradigm proposed to tackle label scarcity in semi-supervised learning.
@@ 2.1 Problem Statement
Formally, given two KGs, $\mathcal{G}_{1}=\{\mathcal{E}_{1},\mathcal{R}_{1},\mathcal{T}_{1}\}$ and $\mathcal{G}_{2}=\{\mathcal{E}_{2},\mathcal{R}_{2},\mathcal{T}_{2}\}$, the task of entity alignment (EA) aims to discover a set of one-to-one equivalent entity pairs =$\mathcal{I}=\{(e_{i},e_{j})\in\mathcal{E}_{1}\times\mathcal{E}_{2}|\;e_{i}\equiv e_{j}\}$ between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, where $e_{i}\in\mathcal{E}_{1}$, $e_{j}\in\mathcal{E}_{2}$, and $\equiv$ indicates an equivalence relationship between $e_{i}$ and $e_{j}$. In many cases, a small set of equivalent entity pairs $\mathcal{S}\subset\mathcal{I}$, known as prior seed alignments, is provided beforehand and used for training. Apart from the entities included in $\mathcal{S}$, there are two sets of unaligned entities $\mathcal{E}^{U}_{1}\subset\mathcal{E}_{1}$ and $\mathcal{E}^{U}_{2}\subset\mathcal{E}_{2}$ in $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively.
In this work, we address real-world scenarios where seed alignments $\mathcal{S}$ are often scarce due to high labeling costs. We focus on the task of pseudo-labeling-based entity alignment, which aims to leverage both seed alignments and unaligned entity pairs to more effectively train an EA model in a transductive semi-supervised setting. This is achieved by selecting a set of unaligned entity pairs as pseudo-labeled alignments $\mathcal{S}_{t}\subset\mathcal{E}^{U}_{1}\times\mathcal{E}^{U}_{2}$ in each iteration $t$, and using $\mathcal{S}_{t}$ to iteratively augment seed alignments $\mathcal{S}$, i.e., $\mathcal{S}\leftarrow\mathcal{S}\cup\mathcal{S}_{t}$, for subsequent model training.
@@ 3.2 Unified Pseudo-Labeling Strategy
After training the EA model, the learned entity embeddings can be used for alignment inference. However, as seed alignments initially provided for training are often limited, the performance of entity alignment can be suboptimal. Therefore, pseudo-labeling strategies can be designed to select a set of unaligned entity pairs as pseudo-labeled alignments, which are used to augment seed alignments for boosting model training. However, as previously discussed in Section 2.2, a naive strategy inevitably introduces a considerable number of pseudo-labeling errors, leading to confirmation bias. To address this issue, we propose UPL, a unified pseudo-labeling framework that explicitly aims to mitigate conflicted misalignments and one-to-one misalignments.
@@ 3.2.1 OT-based Pseudo-Labeling
$T^{*}=\mathop{\arg\min}\limits_{T}\sum_{e_{i}\in\mathcal{E}^{U}_{1}}C_{e_{i},T(e_{i})}.$ (4)
A critical aspect of the above objective is defining a reliable measure of the transport cost. One might directly use the distances between the learned entity embeddings. However, when only a limited number of seed alignments are available for training, the learned entity embeddings can be uninformative, particularly during the early training stages before the EA model has converged. As a result, using these embeddings to calculate the distances for defining the transport cost can be error-prone. To address this, we resort to rectifying the embedding distance using relational neighborhood matching [63], which complements the training of the EA model, especially during the early stages, for learning better entity embeddings and providing a more reliable cost measure for OT modeling. The principle of distance rectification is to leverage relational contexts within local neighborhoods to help determine the extent to which two entities should be aligned. Intuitively, if two entities $e_{i}\in\mathcal{E}_{1}$ and $e_{j}\in\mathcal{E}_{2}$ share more aligned neighboring entities/relations, the distance between their embeddings should be smaller, indicating a higher likelihood of being aligned to each other. Based on this intuition, the transport cost for OT-based pseudo-labeling is defined as follows:
$C_{e_{i},e_{j}}={\color[rgb]{0,0,0}d(e_{i},e_{j})-\lambda s(e_{i},e_{j})},\;e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2},$ (5)
@@
$s(e_{i},e_{j})=\frac{\sum_{(r,e_{k})\in\mathcal{M}_{(e_{i},e_{j})}}\xi_{1}(r,e_{k})\xi_{2}(r,e_{k})}{|\mathcal{N}_{e}(e_{i})|+|\mathcal{N}_{e}({e}_{j})|},$ (6)
@@
where $\mathcal{N}_{e}(e_{i})$ and $\mathcal{N}_{e}(e_{j})$ denote the sets of neighboring entities for $e_{i}$ and $e_{j}$, respectively. $\xi_{1}(r,e_{k})$ and $\xi_{2}(r,e_{k})$ indicate the reciprocal frequency of triplets associated with neighboring tuple $(r,e_{k})$ for $e_{i}$ in $\mathcal{T}_{1}$ and $e_{j}$ in $\mathcal{T}_{2}$, respectively.
@@
The objective in Eq. (4) defines a hard assignment optimization problem, which, however, does not scale well. To enable more efficient optimization and to allow for a more flexible alignment configuration, we reformulate this objective as a discrete OT problem, where the optimal transport plan is considered as a coupling matrix $\bm{P}^{*}\in\mathbb{R}_{+}^{|\mathcal{E}^{U}_{1}|\times|\mathcal{E}^{U}_{2}|}$ between two discrete distributions. Denote $\mu$ and $\nu$ as two discrete probability distributions over all entities $\{e_{i}|e_{i}\in\mathcal{E}^{U}_{1}\}$ and $\{e_{j}|e_{j}\in\mathcal{E}^{U}_{2}\}$, respectively. Without any alignment preference, the two discrete distributions $\mu$ and $\nu$ are assumed to follow a uniform distribution such that $\mu=\frac{1}{|\mathcal{E}^{U}_{1}|}\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\delta_{e_{i}}$ and $\nu=\frac{1}{|\mathcal{E}^{U}_{2}|}\sum_{e_{j}\in\mathcal{E}^{U}_{2}}\delta_{e_{j}}$, where $\delta_{e_{i}}$ and $\delta_{e_{j}}$ are the Dirac function centered on $e_{i}$ and $e_{j}$, respectively. Both $\mu$ and $\nu$ are bounded to sum up to one: $\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\mu(e_{i})=\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\frac{1}{|\mathcal{E}^{U}_{1}|}=1$ and $\sum_{e_{j}\in\mathcal{E}^{U}_{2}}\nu(e_{j})=\sum_{e_{j}\in\mathcal{E}^{U}_{2}}\frac{1}{|\mathcal{E}^{U}_{2}|}=1$. Accordingly, the OT objective is formulated to find the optimal coupling matrix $\bm{P}^{*}$ between $\mu$ and $\nu$:
$\bm{P}^{*}=$ $\mathop{\arg\min}\limits_{\bm{P}\in\Pi(\mu,\nu)}\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\sum_{e_{j}\in\mathcal{E}^{U}_{2}}P_{e_{i},e_{j}}\cdot C_{e_{i},e_{j}},$ (7)
subject to: $\sum_{e_{j}\in\mathcal{E}^{U}_{2}}P_{e_{i},e_{j}}=\mu(e_{i})=\frac{1}{|\mathcal{E}^{U}_{1}|},$
$\sum_{e_{i}\in\mathcal{E}^{U}_{1}}P_{e_{i},e_{j}}=\nu(e_{j})=\frac{1}{|\mathcal{E}^{U}_{2}|},$
$P_{e_{i},e_{j}}\geq 0,\forall e_{i}\in\mathcal{E}^{U}_{1},\forall e_{j}\in\mathcal{E}^{U}_{2},$
@@
where $\Pi(\mu,\nu)=\{\bm{P}\in\mathbb{R}_{+}^{|\mathcal{E}^{U}_{1}|\times|\mathcal{E}^{U}_{2}|}|\;\bm{P}\mathbf{1}_{|\mathcal{E}^{U}_{2}|}=\mu,\;\bm{P}^{\top}\mathbf{1}_{|\mathcal{E}^{U}_{1}|}=\nu\}$ is the set of all joint probability distributions with marginal probabilities $\mu$ and $\nu$, $\mathbf{1}_{n}$ denotes an $n$-dimensional vector of ones. $\bm{P}$ is a coupling matrix signifying probabilistic alignments between two unaligned entity sets $\mathcal{E}^{U}_{1}$ and $\mathcal{E}^{U}_{2}$. Therefore, $P_{e_{i},e_{j}}$ indicates the amount of probability mass transported from $\mu(e_{i})$ to $\nu(e_{j})$. A larger value of $P_{e_{i},e_{j}}$ indicates a higher likelihood of $e_{i}$ and $e_{j}$ being aligned to each other.
@@
$\bm{P}^{*}=\mathop{\arg\min}\limits_{P\in\Pi(\mu,\nu)}\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\sum_{e_{j}\in\mathcal{E}^{U}_{2}}P_{e_{i},e_{j}}\cdot C_{e_{i},e_{j}}+\beta\sum_{e_{i}\in\mathcal{E}^{U}_{1}}\sum_{e_{j}\in\mathcal{E}^{U}_{2}}P_{e_{i},e_{j}}\text{log}P_{e_{i},e_{j}},$ (8)
@@
${\color[rgb]{0,0,0}\widehat{\mathcal{S}}_{t}}=\{(e_{i},e_{j})|\;P^{*}_{e_{i},e_{j}}>\frac{1}{2\cdot\text{min}(|\mathcal{E}^{U}_{1}|,|\mathcal{E}^{U}_{2}|)},\ e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2}\}.$ (9)
@@
Any pseudo-labeled alignment $(e_{i},e_{j}),e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2}$ that satisfies the condition $P^{*}_{e_{i},e_{j}}>\frac{1}{2\cdot\text{min}(|\mathcal{E}^{U}_{1}|,|\mathcal{E}^{U}_{2}|)}$ warrants one-to-one correspondence, such that no conflicted entity pairs, $\{(e_{i},e_{k})|\;e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}\}$ and $\{(e_{l},e_{j})|\;e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}\}$, are selected as pseudo-labeled alignments.
@@
Given the optimized coupling matrix $\bm{P}^{*}\in\mathbb{R}_{+}^{|\mathcal{E}^{U}_{1}|\times|\mathcal{E}^{U}_{2}|}$ subject to the constraints of $\sum_{e_{j}\in\mathcal{E}^{U}_{2}}P^{*}_{e_{i},e_{j}}=\frac{1}{|\mathcal{E}^{U}_{1}|}$ for all rows ($\forall e_{i}\in\mathcal{E}^{U}_{1}$) and $\sum_{e_{i}\in\mathcal{E}^{U}_{1}}P^{*}_{e_{i},e_{j}}=\frac{1}{|\mathcal{E}^{U}_{2}|}$ for all columns ($\forall e_{j}\in\mathcal{E}^{U}_{2}$). Assume that $|\mathcal{E}^{U}_{1}|<|\mathcal{E}^{U}_{2}|$, the decision threshold is $\frac{1}{2\cdot\text{min}(|\mathcal{E}^{U}_{1}|,|\mathcal{E}^{U}_{2}|)}=\frac{1}{2|\mathcal{E}^{U}_{1}|}$. Entity pairs $\{(e_{i},e_{j})|\;P^{*}_{e_{i},e_{j}}>\frac{1}{2|\mathcal{E}^{U}_{1}|},e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2}\}$ are selected as pseudo-labeled alignments.
@@
For each pseudo-labeled alignment $(e_{i},e_{j})$ with a probability value $P^{*}_{e_{i},e_{j}}>\frac{1}{2|\mathcal{E}^{U}_{1}|}$, we can prove that no conflicted entity pairs $\{(e_{i},e_{k})|\;e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}\}$ associated with $e_{i}$ are selected as pseudo-labeled alignments:
$P^{*}_{e_{i},e_{j}}$ $>\frac{1}{2|\mathcal{E}^{U}_{1}|},$
$\sum_{e_{j}\in\mathcal{E}^{U}_{2}}P^{*}_{e_{i},e_{j}}-P^{*}_{e_{i},e_{j}}$ $<\sum_{e_{j}\in\mathcal{E}^{U}_{2}}P^{*}_{e_{i},e_{j}}-\frac{1}{2|\mathcal{E}^{U}_{1}|},$
$\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}P^{*}_{e_{i},e_{k}}+P^{*}_{e_{i},e_{j}}-P^{*}_{e_{i},e_{j}}$ $<\frac{1}{|\mathcal{E}^{U}_{1}|}-\frac{1}{2|\mathcal{E}^{U}_{1}|},$
$\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}P^{*}_{e_{i},e_{k}}$ $<\frac{1}{2|\mathcal{E}^{U}_{1}|}.$ (10)